# RIME: Robust Preference-based Reinforcement Learning with Noisy Preferences

### v3

@@ 5.3 Ablation Study
Table 4: Results of different robust training methods on tasks from Meta-world and DMControl with “mistake” teacher and error rate as $\epsilon=0.3$. The result shows the mean and standard deviation of the metric averaged over 5 runs.
@@ Acknowledgments
We are grateful to the anonymous reviewers and editors for their insightful suggestions. This work was partially supported by the National Natural Science Foundation of China under Grants 62271485, U1909204 and 62303462, the Provincial Key Research and Development Program of Zhejiang (project number: 2022C01129), Ningbo International Science and Technology Cooperation Project (2023H020), and Beijing Natural Science Foundation under Grant L233005. Lastly, special thanks go to Tongtian Yue, Ruixi Qiao, and Yueyang Yang for their assistance in data annotation, engaging discussions, and meticulous review of draft versions.
@@ Appendix B Effects of biased reward model
Consider a Markov Decision Process characterized by the state transition function $P$, ground-truth reward function $r^{*}$, and discount factor $\gamma$. Let $Q^{\pi}$ denote the Q-function for policy $\pi$ with respect to the learned reward model $\hat{r}_{\psi}$. Then the error between $Q^{\pi}$ and the optimal Q-function $Q^{*}$ is upper-bounded by the fitting error $\delta$ of the reward model:
$\operatorname*{\mathbb{E}}_{(\mathbf{s},\mathbf{a})\sim\rho^{\pi}}\left|Q^{\pi}(\mathbf{s},\mathbf{a})-Q^{*}(\mathbf{s},\mathbf{a})\right|\leq\frac{\delta}{1-\gamma\delta}$ (12)
@@
$\operatorname*{\mathbb{E}}_{(\mathbf{s},\mathbf{a})\sim\rho^{\pi}}\left|Q^{\pi}(\mathbf{s},\mathbf{a})-Q^{*}(\mathbf{s},\mathbf{a})\right|$ (13)
$=$ $\operatorname*{\mathbb{E}}_{(\mathbf{s},\mathbf{a})\sim\rho^{\pi}}\left|\hat{r}_{\psi}(\mathbf{s},\mathbf{a})+\gamma\sum_{\mathbf{s}^{\prime}}P(\mathbf{s}^{\prime}|\mathbf{s},\mathbf{a})\sum_{\mathbf{a}^{\prime}}\pi(\mathbf{a}^{\prime}|\mathbf{s}^{\prime})Q^{\pi}(\mathbf{s}^{\prime},\mathbf{a}^{\prime})-r^{*}(\mathbf{s},\mathbf{a})-\gamma\sum_{\mathbf{s}^{\prime}}P(\mathbf{s}^{\prime}|\mathbf{s},\mathbf{a})\max_{\mathbf{a}^{\prime}}Q^{*}(\mathbf{s}^{\prime},\mathbf{a}^{\prime})\right|$ (14)
$\leq$ $\operatorname*{\mathbb{E}}_{(\mathbf{s},\mathbf{a})\sim\rho^{\pi}}\left|\hat{r}_{\psi}(\mathbf{s},\mathbf{a})-r^{*}(\mathbf{s},\mathbf{a})\right|+\gamma\operatorname*{\mathbb{E}}_{(\mathbf{s},\mathbf{a})\sim\rho^{\pi}}\sum_{\mathbf{s}^{\prime}}P(\mathbf{s}^{\prime}|\mathbf{s},\mathbf{a})\left|\sum_{\mathbf{a}^{\prime}}\pi(\mathbf{a}^{\prime}|\mathbf{s}^{\prime})Q^{\pi}(\mathbf{s}^{\prime},\mathbf{a}^{\prime})-\max_{\mathbf{a}^{\prime}}Q^{*}(\mathbf{s}^{\prime},\mathbf{a}^{\prime})\right|$ (15)
$=$ $\delta+\gamma\operatorname*{\mathbb{E}}_{(\mathbf{s},\mathbf{a})\sim\rho^{\pi}}\sum_{\mathbf{s}^{\prime}}P(\mathbf{s}^{\prime}|\mathbf{s},\mathbf{a})\left|\sum_{\mathbf{a}^{\prime}}\pi(\mathbf{a}^{\prime}|\mathbf{s}^{\prime})(Q^{\pi}(\mathbf{s}^{\prime},\mathbf{a}^{\prime})-\max_{\mathbf{a}^{\prime\prime}}Q^{*}(\mathbf{s}^{\prime},\mathbf{a}^{\prime\prime}))\right|$ (16)
$\leq$ $\delta+\gamma\operatorname*{\mathbb{E}}_{(\mathbf{s},\mathbf{a})\sim\rho^{\pi}}\sum_{\mathbf{s}^{\prime}}P(\mathbf{s}^{\prime}|\mathbf{s},\mathbf{a})\sum_{\mathbf{a}^{\prime}}\pi(\mathbf{a}^{\prime}|\mathbf{s}^{\prime})(\max_{\mathbf{a}^{\prime\prime}}Q^{*}(\mathbf{s}^{\prime},\mathbf{a}^{\prime\prime})-Q^{\pi}(\mathbf{s}^{\prime},\mathbf{a}^{\prime}))$ (17)
$\leq$ $\delta+\gamma\operatorname*{\mathbb{E}}_{(\mathbf{s},\mathbf{a})\sim\rho^{\pi}}\sum_{\mathbf{s}^{\prime}}P(\mathbf{s}^{\prime}|\mathbf{s},\mathbf{a})\sum_{\mathbf{a}^{\prime}}\pi(\mathbf{a}^{\prime}|\mathbf{s}^{\prime})(Q^{*}(\mathbf{s}^{\prime},\mathbf{a}^{\prime})-Q^{\pi}(\mathbf{s}^{\prime},\mathbf{a}^{\prime}))$ (18)
$=$ $\delta+\gamma\operatorname*{\mathbb{E}}_{(\mathbf{s}^{\prime},\mathbf{a}^{\prime})\sim\rho^{\pi}}\left|Q^{\pi}(\mathbf{s}^{\prime},\mathbf{a}^{\prime})-Q^{*}(\mathbf{s}^{\prime},\mathbf{a}^{\prime})\right|$ (19)
$\leq$ $\delta+\gamma\delta^{2}+\gamma^{2}\delta^{3}+\dots$ (20)
$\leq$ $\frac{\delta}{1-\gamma\delta}$ (21)

### v4

@@ 5.3 Ablation Study
Table 4: Results of different robust training methods with “mistake” teacher and error rate as $\epsilon=0.3$.
@@ Acknowledgments
We are grateful to the anonymous reviewers and editors for their insightful suggestions. This work was partially supported by the National Science and Technology Major Project (2022ZD0117102), the National Natural Science Foundation of China under Grants 62271485, U1909204 and 62303462, the Provincial Key Research and Development Program of Zhejiang (project number: 2022C01129), Ningbo International Science and Technology Cooperation Project (2023H020), and Beijing Natural Science Foundation under Grant L233005. Lastly, special thanks go to Tongtian Yue, Ruixi Qiao, and Yueyang Yang for their assistance in data annotation, engaging discussions, and meticulous review of draft versions.
@@ Appendix B Effects of biased reward model
Consider a Markov Decision Process characterized by the state transition function $P$, ground-truth reward function $r^{*}$, and discount factor $\gamma$. Let $Q^{\pi}_{\psi}$ and $Q^{\pi}_{*}$ denote the Q-function for policy $\pi$ with respect to the learned reward model $\hat{r}_{\psi}$ and ground-truth reward function $r^{*}$, respectively. Then the error of policy estimation between $Q^{\pi}_{\psi}$ and $Q^{\pi}_{*}$ is upper-bounded by the fitting error $\delta$ of the reward model:
$\operatorname*{\mathbb{E}}_{(\mathbf{s},\mathbf{a})\sim\rho^{\pi}}\left|Q^{\pi}_{\psi}(\mathbf{s},\mathbf{a})-Q^{\pi}_{*}(\mathbf{s},\mathbf{a})\right|\leq\frac{\delta}{1-\gamma}$ (12)
@@
$\operatorname*{\mathbb{E}}_{(\mathbf{s},\mathbf{a})\sim\rho^{\pi}}\left|Q^{\pi}_{\psi}(\mathbf{s},\mathbf{a})-Q^{\pi}_{*}(\mathbf{s},\mathbf{a})\right|$
$=$ $\operatorname*{\mathbb{E}}_{(\mathbf{s},\mathbf{a})\sim\rho^{\pi}}\left|\hat{r}_{\psi}(\mathbf{s},\mathbf{a})+\gamma\sum_{\mathbf{s}^{\prime}}P(\mathbf{s}^{\prime}|\mathbf{s},\mathbf{a})\sum_{\mathbf{a}^{\prime}}\pi(\mathbf{a}^{\prime}|\mathbf{s}^{\prime})Q^{\pi}_{\psi}(\mathbf{s}^{\prime},\mathbf{a}^{\prime})-r^{*}(\mathbf{s},\mathbf{a})-\gamma\sum_{\mathbf{s}^{\prime}}P(\mathbf{s}^{\prime}|\mathbf{s},\mathbf{a})\sum_{\mathbf{a}^{\prime}}\pi(\mathbf{a}^{\prime}|\mathbf{s}^{\prime})Q^{\pi}_{*}(\mathbf{s}^{\prime},\mathbf{a}^{\prime})\right|$
$\leq$ $\operatorname*{\mathbb{E}}_{(\mathbf{s},\mathbf{a})\sim\rho^{\pi}}\left|\hat{r}_{\psi}(\mathbf{s},\mathbf{a})-r^{*}(\mathbf{s},\mathbf{a})\right|+\gamma\operatorname*{\mathbb{E}}_{(\mathbf{s},\mathbf{a})\sim\rho^{\pi}}\sum_{\mathbf{s}^{\prime}}P(\mathbf{s}^{\prime}|\mathbf{s},\mathbf{a})\sum_{\mathbf{a}^{\prime}}\pi(\mathbf{a}^{\prime}|\mathbf{s}^{\prime})\left|Q^{\pi}_{\psi}(\mathbf{s}^{\prime},\mathbf{a}^{\prime})-Q^{\pi}_{*}(\mathbf{s}^{\prime},\mathbf{a}^{\prime})\right|$
$=$ $\delta+\gamma\operatorname*{\mathbb{E}}_{(\mathbf{s}^{\prime},\mathbf{a}^{\prime})\sim\rho^{\pi}}\left|Q^{\pi}_{\psi}(\mathbf{s}^{\prime},\mathbf{a}^{\prime})-Q^{\pi}_{*}(\mathbf{s}^{\prime},\mathbf{a}^{\prime})\right|$
$\leq$ $\delta+\gamma\delta+\gamma^{2}\delta+\gamma^{3}\delta+\dots$
$\leq$ $\frac{\delta}{1-\gamma}$ (13)